OpenStudy (anonymous):

whats the intersection points for 3y+6x=3 and x-y=-7?

OpenStudy (unklerhaukus):

if you add y to both sides of the second equation, you'll get x= ... sub this into the first equation and solve for y

OpenStudy (anonymous):

still confused

OpenStudy (unklerhaukus):

x-y = -7 adding y to both sides of the second equation x-y +y = -7 +y simplfying x = -7 +y

OpenStudy (midhun.madhu1987):

 just add these 2 equations.. and y will be eliminated... find x,